Operating Systems and Systems Integration

Quiz 11: Secure Shell

Do not refer to any printed material. No talking or discussion is allowed until the answer sheets are all collected.

Review:

1. Secure Shell is:

(A) what Nick uses to connect to his desktop machine to show the lectures and laboratory work (B) not useful for remote administration (C) always requires you to use a special hardware password device to authenticate yourself, rather

like a pocket calculator (D) is just another name for bash (E) is not useful for port forwarding, or carrying other network traffic, and does not use encryption

2. If you cannot connect to a remote computer using secure shell (ssh), you should: (A) Call the police (B) Check whether the secure shell server is running on the remote computer, because an ssh

client needs a server to connect to (C) Shut down your own computer immediately and check for security violations (D) Check whether your computer has aids (Acquired Immune Deficiency Syndrome) (E) Use the \underline{S}uper \underline{S}oftware \underline{H}ickup remover to get your SSH working properly again.

3. The ssh agent is:

(A) A Cyber James Bond

(B) A program that holds the decrypted private key

(C) A tool for attacking other computers

(D) A cracking tool

(E) used to kill people. Preparation:

4. ssh uses keys. The host keys are used to:

(A) authenticate the server and are used to begin the encryption

(B) authenticate the user

(C) block data transmission

(D) prevent any files being transferred

(E) terminate the agent

5. The user public and private keys are used to:

(A) authenticate the server and are used to begin the encryption

(B) authenticate the user

(C) block data transmission

(D) prevent any files being transferred

(E) terminate the agent

Nick Urbanik <nicku(at)nicku.org>  ver. 1.0